OpenStudy (anonymous):

A 1020-kg car is on an inclined plane inclined at an angle of 30 degrees. A cable is connected to the car, stretched over a pulley at the end of the incline and connected to a 11900-N counterweight (a) Determine the acceleration (in m/s/s) of the car. If the car accelerates down the inclined plane, then enter a negative answer. Ignore any effects of friction. m/s/s (b) Determine the tension (in Newtons) in the cable. N (c) What mass (in kg) should the counterweight have in order for the car to move down the incline at an acceleration of 2.43 m/s/s?

OpenStudy (anonymous):

ok what's the question?

OpenStudy (anonymous):

its up there

OpenStudy (anonymous):

sorry internet acting up now I see it

OpenStudy (anonymous):

ok first draw a free body diagram of both the counterweight and the car

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

from this we can get an equation for the counterweight being\[W-T=m _{2}\]

OpenStudy (anonymous):

correction \[W-T=m _{2}a \]

OpenStudy (anonymous):

Now for the car we have\[T-Wcos \theta=m _{1}a\]

OpenStudy (anonymous):

now we can add the two together to get\[W _{2}-W _{1}\cos \theta=m ^{1}a+m _{2}a\]

OpenStudy (anonymous):

\[\frac{ W _{2}-W _{1}\cos \theta }{ m _{1}+m _{2} }=a\]

OpenStudy (anonymous):

now you can plug in your given amounts and solve for a

OpenStudy (anonymous):

once you solve for a you can then plug that in to find T

OpenStudy (michele_laino):

Sorry @recon14193 what you got for acceleration value, nameli option a), please?

OpenStudy (anonymous):

a is 3.09

OpenStudy (anonymous):

i cant figure out T

OpenStudy (michele_laino):

@chicagogirl5673 I got a=3.38 m/sec^2